Give the formulas for and plot AFC,MC,AVC and AC if the cost function is 1.C=10+10Q, 2. C=10+q2, 3.C=10+10q-4q2 +q3.


1
Expert's answer
2021-10-24T18:10:56-0400

"AFC = \frac{FC}{Q}"

MC is the derivative of C with respect to q.

"AVC =\frac{ VC}{Q}"

"AC = \frac{C}{Q}"

The FC is anything in the C equation that isn't attached to a Q. The VC is anything that is attached to a Q.

a.

a.

C = 10 + 10Q

FC = 10

"AFC = \frac{10}{Q}"

MC = 10

AVC = 10

"AC = \frac{10}{Q} + 10"


b.

"C = 10 + Q^2"

FC = 10

"AFC =\frac{10}{Q}"

MC = 2Q

AVC = Q

"AC = \frac{10}{Q} + Q"


c.

"C = 10 + 10Q - 2Q^2 + Q^3"

FC = 10

"AFC = \frac{10}{Q}"

"MC = 10 - 4Q + 3Q^2"

"AVC = 10 - 2Q + Q^2"

"AC = \frac{10}{Q} + 10 - 2Q + Q^2"


Graphs

a.






b.






c.